OpenStudy (anonymous):

what is the derivative of sin^2 x??

OpenStudy (reemii):

it is a compound function. use \((f(g(x)))' = f'(g(x))g'(x)\). here, \(f(x)=x^2\), and \(g(x)=\sin x\). therefore, \(((g)^2)' =( 2g) \times g'(x)\).

OpenStudy (zzr0ck3r):

ok like @reemi said you have a composite function (sinx)^2 first we deal with the outside function which is something^2 so we take the derivative of the outside 2(something) and multiply this by the derivative of that something 2sin^(x)cos(x)

OpenStudy (zzr0ck3r):

its the same thing as (2x+3)^2 accept here instead of 2x+3 you have sin(x)

OpenStudy (zzr0ck3r):

please don't tell me to stop typing then delete the comment....

OpenStudy (zzr0ck3r):

so back to the math. so we have 2sinxcosx now if you would like to apply some trig identity to get it to look nicer you should have one of the form @jack17 gave.

OpenStudy (anonymous):

2sinxcosx @jessie096

OpenStudy (zzr0ck3r):

dude makes me like you:P

OpenStudy (zzr0ck3r):

ah nm he deleted it all

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Wanna help me with a math question

OpenStudy (zzr0ck3r):

lol oh god. is it the integral to find the volume of water on earth?

OpenStudy (zzr0ck3r):

what class is it for?

OpenStudy (anonymous):

It's a really easy question :)

OpenStudy (zzr0ck3r):

lol

OpenStudy (anonymous):

But seriously, it is

OpenStudy (zzr0ck3r):

let me go outside and smoke real fast, paste it to me.

OpenStudy (zzr0ck3r):

I think you have a much better grip on non analysis calculus, so if you cant get it....

OpenStudy (anonymous):

Let W be the subspace of vectors in R^3 that are orthogonal to w = (-3,5,1). Find a basis for W. What is the dimension of the subspace? Give a geometrical description.

OpenStudy (zzr0ck3r):

I don't remember all methods from calculus

OpenStudy (zzr0ck3r):

ahh I can do this.

OpenStudy (zzr0ck3r):

brb

OpenStudy (anonymous):

lol everyone just have a conversation on my question :3 anyways thanks!!!

OpenStudy (reemii):

so the word is "composite" not "compound" ?